OpenStudy (anonymous):

I have just started algebra 1 honors and I need help. Would someone work this out step by step PLEASE 6 + v/-8 = 4/7

OpenStudy (anonymous):

You want to isolate the \(v\).

OpenStudy (anonymous):

The easiest way to do this is to do the order of operations backwards.

OpenStudy (anonymous):

This is because if you multiply and there is still addition remaining, you have to distribute it, and that makes things messy.

OpenStudy (anonymous):

So the first step is to subtract

OpenStudy (anonymous):

To isolate \(v\) you need \(6\) on the other side. \[ \color{red}6+\frac v{-8}=\frac 47 \]

OpenStudy (anonymous):

To do this we need to subtract \(6\) from both sides.

OpenStudy (anonymous):

The result is \[ 6-6+\frac{v}{-8}=\frac 47-6 \]However the \(6\) cancels out by design and we get: \[ \frac{v}{-8}=\frac 47 -6 \]

OpenStudy (anonymous):

The next step is to simplify the right side: \[ \frac 47 -6 \]

OpenStudy (anonymous):

We need the denominator to be the same, so we change \(6\): \[ 6 = \frac 61 = \frac 61 \times \frac 77=\frac {42}7 \]

OpenStudy (anonymous):

Now we can simplify: \[ \frac 47 - \frac {42}7=\frac{4-42}7=\frac{-38}7 \]

OpenStudy (anonymous):

Our equation simplified is: \[ \frac{v}{-8}=\frac{-38}{7} \]

OpenStudy (anonymous):

Now let me know if anything confused you so far.

OpenStudy (anonymous):

I'm sorry, where did 7/7 come from?

OpenStudy (anonymous):

Well, \(7/7 = 1\). So we multiplied \(6\) by \(1\).

OpenStudy (anonymous):

This didn't change anything because anything times \(1\) is itself.

OpenStudy (anonymous):

Okay, so you know how \(12/3=4\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well we are doing something similar, but we are going backwards.

OpenStudy (anonymous):

We are trying to find a number where: \(??/7=6\)

OpenStudy (anonymous):

Because we need \(7\) in the denominator (bottom part) in order to add the fractions.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

No, sorry

OpenStudy (anonymous):

Do you know how to add fractions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is: \[ \frac 45 + \frac 28 \]

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

4/5 + 2/8 = + 10/40 +42/40 which equals 1 2/40 or 1 1/20

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oops, missed a number 32/40 + 10/40 = 42/40 then reduce to simple form

OpenStudy (anonymous):

Okay, good.

OpenStudy (anonymous):

What is \(\frac 47 - 6\)?

OpenStudy (anonymous):

4/7 - 6/1 = 4/7 - 42/7 = -38/7

OpenStudy (anonymous):

okay, well if you don't understand what I did, then just understand that: \[ \frac v{-8}=\frac 47-6 \]Simplified to :\[ \frac v{-8}=\frac{-38}7 \]Okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then, to isolate the \(v\), what is left to do?

OpenStudy (anonymous):

You need to get rid of \(-8\), so how do you do it?

OpenStudy (anonymous):

subtract 1/8 from both sides?

OpenStudy (anonymous):

Nope. subtraction is the opposite of addition. We're dividing. We need the opposite of division.

OpenStudy (anonymous):

multiply 1/8 on both sides

OpenStudy (anonymous):

Close, but: \[ \frac{v}{-8}\times \frac 18 = \frac{v}{-64} \]

OpenStudy (anonymous):

So we multiply, but multiply by something else.

OpenStudy (anonymous):

We want to get \(v\) alone.

OpenStudy (anonymous):

I'm stuck

OpenStudy (anonymous):

We want to multiply by -8

OpenStudy (anonymous):

so would it be -8/1

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

\[ v=\frac{-38}{7}\times \frac{-8}1 \]

OpenStudy (anonymous):

so now we get the denominators to match?

OpenStudy (anonymous):

-38/7 x -8/1 = -38/7 x -56/7 = 2128/49 = 43.45 (reduced) ???

OpenStudy (anonymous):

Wio, are you still here?

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

What you did is correct.

OpenStudy (anonymous):

However, you don't need to make denominators equal when you multiply fractions, so you didn't more work than necessary. it is still correct though.

OpenStudy (anonymous):

so when I multiply, I can cross multiple and be done with it?

OpenStudy (anonymous):

Wio, THANK YOU SO MUCH!!!!!!!

OpenStudy (anonymous):

You found the answer \(v\approx 43.4\)

OpenStudy (anonymous):

Thank you again! Have a great night